OpenStudy (anonymous):

differentiate and simplify csc^2 theta + cot^2 theta

OpenStudy (anonymous):

Remember that \[\csc ^{2}\] is just =\[-\cot ^{2}(\theta)\csc ^{2}(\theta)\]

OpenStudy (anonymous):

got that part. then what?

OpenStudy (anonymous):

Let's say that theta = x, also note that 1+cot^2x=csc^2x d/dx[csc^2x-cot^2x] = = -cot^xcsc^2x-csc^2x = = -csc^2x(cot^2x+1) = = csc^x(csc^2x) = = -csc^4x the answer is -csc^4(theta)

OpenStudy (anonymous):

thanks!